# Acceleration of Heavy Ions at Non-Relativistic Collisionless Shocks

### v1

@@ 2 Simulations
We run simulations with Alfvénic Mach numbers of $M=\{5,20,40\}$. The Mach number, $M=M_{A}\simeq M_{s}$, is defined as the ratio of the upstream bulk flow speed $v_{sh}$ (measured in the downstream rest frame) to the Alfvén speed $v_{A}$ (for the Alfvénic Mach number $M_{A}$) or to the sound speed (for the sonic Mach number $M_{s}$). We also define the quantity $E_{sh}=mv_{sh}^{2}/2$ to which energies are normalized. In all simulations, we use 25 particles per cell for H and 4 particles per cell for heavier species, with a resolution of 2 cells per $d_{i}$. The shock evolution is followed for hundreds of ion cyclotron times until the expected power law distribution is established.
@@ 3 High-$M$ regime
We first consider a benchmark run representative of strong shocks with $M=40$. Figure 1a shows the downstream energy spectra for different species at $t=500\,\omega_{c}^{-1}$, measured over an integration window of $450\,d_{i}$ downstream of the shock. This choice ensures that the spectra reflect only the latest acceleration stages, without contamination from earlier times, when injection of heavy ions has not started, yet. The results are compared with a run using the same parameters but including only H (dashed lines). When heavy ions are included, the normalization of the H non-thermal tail decreases and ends up below the He tail, which dominates energetically; the CNO ions remain subdominant, with spectra about an order of magnitude below both H and He, according to the expected enhancement $\propto(A/Q)^{2}$. In addition, H and He particles reach maximum energies $\sim 2$ times higher than H in the H-only run. The modification of the H spectrum and the higher maximum energies attained indicate that heavy ions play a dynamically important role in the acceleration process, as we will discuss in the next Section. Finally, in the presence of heavy elements, the peaks of the He and CNO Maxwellians scale correctly with $A$, consistent with the test-particle results of CSY17, whereas the H peak does not, a result of the extra energy that the shock channels into heavy ions, which is subtracted from the heating of thermal H.
@@ 4 The scaling with $M$
Figure 2 shows the ion enhancement as a function of $M$ for both He and CNO. For weaker, lower-Mach number shocks, heavy ions experience little to no enhancement in injection. However, at stronger shocks $K_{sp}$ becomes significant and increases with $M$. For instance, at $M=40$, He is enhanced by a factor of approximately 30, while CNO ions show an enhancement of about 250. It is worth noting that when heavy ions are treated as test particles, the ion enhancement saturates for $M\gtrsim 10$ (see Figure 2 in CYS17). In contrast, our results show a continued increase with $M$, although the growth is less dramatic beyond $M=20$ compared to the steep rise observed from $M=5$ to $M=20$. This difference, however, does not alter the $(A/Q)^{2}$ scaling reported in CYS17, which successfully reproduces the observed elemental abundances in Galactic CRs.
@@ 5.2 Overall spectra of He and heavy ions
5.3 Post-shock ion temperatures
Another potential observable stemming from our results is that the post-shock temperatures of different species may not simply scale with $A$, especially with H falling out of the correlation, as discussed above. Models of X-ray bremsstrahlung or Balmer emission that rely on inferring the H temperature from Rankine–Hugoniot conditions or assume that all the species are in thermal equilibrium may lead to inaccurate interpretations.
@@ 6 Conclusion
One final comment is that in this work we focused on quasi-parallel shocks, only. Though oblique and quasi-perpendicular configurations hinder the spontaneous injection of thermal particles, it was recently found that high-$M$ shocks may indeed accelerate particles at non-negligible levels (L. Orusa & D. Caprioli, 2023; L. Orusa & V. Valenzuela-Villaseca, 2025; L. Orusa et al., 2025). Since this requires full 3D simulations and is thus computationally very expensive, we defer such an investigation to future works.
We would like to thank Marwah Roussi and Elyssa Brooks for having worked on preliminary runs of shocks including heavy ions. This research was partially supported by NSF grants AST-2510951 and AST-2308021 and by NASA grant 80NSSC18K1726. L.O. acknowledges the support of the Multimessenger Plasma Physics Center (MPPC), NSF grant PHY2206607. We would like to thank The University of Chicago Research Computing Center for providing the computational resources to conduct this research.

### v2

@@ 2 Simulations
We run simulations with Alfvénic Mach numbers of $M=\{5,20,40\}$. The Mach number, $M=M_{A}\simeq M_{s}$, is defined as the ratio of the upstream bulk flow speed $v_{sh}$ (measured in the downstream rest frame) to the Alfvén speed $v_{A}$ (for the Alfvénic Mach number $M_{A}$) or to the sound speed (for the sonic Mach number $M_{s}$). $c$ is chosen for the different simulations such that $v_{sh}=0.1c$. We also define the quantity $E_{sh}=mv_{sh}^{2}/2$ to which energies are normalized. In all simulations, we use 25 particles per cell for H and 4 particles per cell for heavier species, with a resolution of 2 cells per $d_{i}$. The shock evolution is followed for hundreds of ion cyclotron times until the expected power law distribution is established.
@@ 3 High-$M$ regime
We first consider a benchmark run representative of strong shocks with $M=40$. Figure 1a shows the downstream energy spectra for different species at $t=500\,\omega_{c}^{-1}$, measured over an integration window of $450\,d_{i}$ downstream of the shock. This choice ensures that the spectra reflect only the latest acceleration stages, without contamination from earlier times, when injection of heavy ions has not started, yet. The results are compared with a run using the same parameters but including only H (dashed lines). When heavy ions are included, the normalization of the H non-thermal tail decreases and ends up below the He tail, which dominates energetically; the CNO ions remain subdominant, with spectra about an order of magnitude below both H and He, according to the expected enhancement $\propto(A/Q)^{2}$. In addition, H and He particles reach maximum energies $\sim 2$ times higher than H in the H-only run. The modification of the H spectrum and the higher maximum energies attained indicate that heavy ions play a dynamically important role in the acceleration process, as we will discuss in the next Section. Finally, in the presence of heavy elements, the peaks of the He and CNO Maxwellians scale with $A$ as reported by CSY17, whereas the H peak does not, a result of the extra energy that the shock channels into heavy ions, which is subtracted from the heating of thermal H. Models of X-ray bremsstrahlung or Balmer emission that rely on inferring the H temperature from Rankine–Hugoniot conditions or assume that all the species are in thermal equilibrium may be affected by this effect.
@@ 4 The scaling with $M$
Figure 2 shows the ion enhancement as a function of $M$ for both He and CNO. For weaker, lower-Mach number shocks, heavy ions experience little to no enhancement in injection. However, at stronger shocks $K_{sp}$ becomes significant and increases with $M$. For instance, at $M=40$, He is enhanced by a factor of approximately 30, while CNO ions show an enhancement of about 250. It is worth noting that when heavy ions are treated as test particles, the ion enhancement saturates for $M\gtrsim 10$ (see figure 2 in CYS17). In contrast, our results show a continued increase with $M$, although the growth is less dramatic beyond $M=20$ compared to the steep rise observed from $M=5$ to $M=20$. This difference, however, does not alter the $(A/Q)^{2}$ scaling reported in CYS17, which successfully reproduces the observed elemental abundances in Galactic CRs. It is worth mentioning that also the sputtering of dust undergoing DSA may be a promising mechanism for heavy metal enrichment (e.g., J. Meyer et al., 1997; D. C. Ellison et al., 1997; V. Tatischeff et al., 2021).
@@ 6 Conclusion
One final comment is that in this work we focused on quasi-parallel shocks in the warm interstellar medium, only. Different initial abundances and ionization states, such as in solar energetic particles (e.g., A. J. Tylka et al., 2005; D. V. Reames, 2013) or in superbubbles (e.g., J. C. Higdon & R. E. Lingenfelter, 2005; V. Tatischeff et al., 2021) may lead to different scaling of the enhancements with $A/Q$ (M. I. Desai et al., 2016; J.-Y. Lee et al., 2024). Moreover, though oblique and quasi-perpendicular configurations hinder the spontaneous injection of thermal particles, it was recently found that high-$M$ shocks may indeed accelerate particles at non-negligible levels (L. Orusa & D. Caprioli, 2023; L. Orusa & V. Valenzuela-Villaseca, 2025; L. Orusa et al., 2025). Since this requires full 3D simulations and is thus computationally very expensive, we defer such an investigation to future works.
We would like to thank the referee for having contributed to improve the manuscript; Marwah Roussi and Elyssa Brooks for having worked on preliminary runs of shocks including heavy ions; and Luke O’C. Drury and Anatoly Spitkovsky for stimulating discussions and technical/scientific advice. This research was partially supported by NASA grant 80NSSC18K1726, NSF grants AST-2510951 and AST-2308021 to D.C., and by NSF grant PHY-2309135 to the Kavli Institute for Theoretical Physics. L.O. acknowledges the support of the Multimessenger Plasma Physics Center (MPPC), NSF grant PHY2206607. We would like to thank The University of Chicago Research Computing Center for providing the computational resources to conduct this research.